find the power of i 
i18

Found 2 solutions by AnlytcPhil, kim15432:
Answer by AnlytcPhil(1806)

find the power of i 
i18

All you have to know for problems like this is that

i2 = -1

And that

(-1)even power = 1 

and 

(-1)odd power = -1

i18 = i2·9 = (i2)9 = (-1)9 = -1 since 9 is an odd power

That's the answer.

=======================================================

Suppose you had the problem i73, then

i73 = i72+1 = i72i1 = (i2)36i1 = (-1)36i1 = 

(1)i1   <-----(because 36 is an even power)

Erase the 1's

Answer: i

Edwin



Answer by kim15432(2)